OpenStudy (anonymous):

Identify the 12th term of a geometric sequence where a1 = 8 and a6 = -8,192.

OpenStudy (anonymous):

sorry i know they're all the same

OpenStudy (anonymous):

same thing ar^0 + ar^1 a is 8 -8192=8*r^5

OpenStudy (campbell_st):

then using the term in a GP formula \[a _{n} = a _{1}r^{n -1}\] \[-8192 = 8r^5\] solve for r

OpenStudy (anonymous):

hmmmm lets see...

OpenStudy (campbell_st):

then use the formula above with n = 12

OpenStudy (campbell_st):

can you find r..?

OpenStudy (campbell_st):

divide both sides by 8 \[r^5 = -\frac{8192}{8} ..... then...... r^5 = -1024\]

OpenStudy (anonymous):

mk i got that..... but do u do nething with the ^5?

OpenStudy (campbell_st):

\[r = \sqrt[5]{-1024}\]

OpenStudy (anonymous):

mmmmmkkkkkkkkk................................ im sorry im lost

OpenStudy (campbell_st):

well r = -4 so you now have the information you need a = 8 r = -4 and n = 12 substitute it into the formula \[a _{n} = a _{1} r^{n -1}\]

OpenStudy (anonymous):